1. What is the calculated value of the cell potential at 298K for an electrochemical cell with the following reaction, when the Cl2 pressure is 9.21×10-4 atm, the Cl- concentration is 1.38M, and the Pb2+ concentration is 1.19M ?

Cl2(g) + Pb(s) -------> 2Cl-(aq) + Pb2+(aq)

Answer: _______ V

The cell reaction as written above is spontaneous for the concentrations True or False?

(This is all part of one question)
Answer #1

Reaction at anode:Pb (s)   Pb 2+ (aq) + 2 e -

Reaction at cathode: Cl 2 (g) + 2 e -   2 CI -(aq)

Here n= no of electrons transferred = 2

Q = [ Pb 2+ ] [ 2 CI - ] / ( P Cl 2 )

E 0c = E 0 cathode - E 0 anode

= 1.36 V - ( - 0.126 V)

= 1.486 V

Emf of above cell is calculated by using Nernst equation as

E cell = E 0 cell - [ 2.303 R T / n F ] log Q

Where , n is no of electrons transferred in the reaction ,

Q is the reaction quotient.

E cell = 1.486 - 2.303 R T / n F log Q

= 1.486 - 0.0591 / 2 log [ Pb 2+ ] [ CI - ] 2 / ( P Cl 2 )

= 1.486 - 0.02955 log 1.19  ( 1.38 ) 2 / ( 9.21 x 10 -04 )

= 1.486 - 0.02955 x 3.3910

= 1.486 - 0.100

= 1.386 V

Reaction is spontaneous .

ANSWER : TRUE

E cell > 0 and hence G 0 < 0 . So reaction is spontaneous.